OpenStudy (anonymous):

y=6x-1 and y=1/6x-5= perpendicular right? y-3=4(x+2) and y+3=4(x+5)+ parallel right?

OpenStudy (shamim):

ur given 1st 2 line r not perpendicular

OpenStudy (shamim):

slope of ur 1st line is\[m _{1}=6\]

OpenStudy (luigi0210):

Hm? The first line is not perpendicular

OpenStudy (shamim):

slope of ur 2nd line is\[m _{2}=\frac{ 1 }{ 6 }\]

OpenStudy (shamim):

and \[m _{1}m _{2}=6 \times \frac{ 1 }{ 6 }=1\]

OpenStudy (shamim):

it should b \[-1\]for perpendicular

OpenStudy (anonymous):

so neither

OpenStudy (shamim):

ya latest 2 lines r parallel

OpenStudy (shamim):

its true

OpenStudy (shamim):

u can ask any question if u want

OpenStudy (anonymous):

so its perpendicular

OpenStudy (shamim):

actually u asked 2 questions here

OpenStudy (shamim):

u hv 4 equation of line here

OpenStudy (shamim):

ur 1st pair of lines r not perpendicular

OpenStudy (shamim):

bt ur 2nd pair of lines r parallel

OpenStudy (anonymous):

the first pair of lines are neither

OpenStudy (shamim):

ur 1st pair of lines r not parallel n not perpendicular

OpenStudy (shamim):

if\[m _{1}=m _{2}\]then those lines r parallel

OpenStudy (shamim):

and if \[m _{1}m _{2}=-1\]then those lines r perpendicular

OpenStudy (anonymous):

ok, i got it.

OpenStudy (shamim):

ok